OpenStudy (anonymous):

Addison earned a gross income of $42,750 last year. She made $537.56 in IRA contributions, donated $918 to her favorite charity and paid $1,351.35 in home mortgage interest. If Addison claims a standard deduction of $5,700 and her exemption is $3,650, what is her taxable income?

OpenStudy (confluxepic):

\[\huge \color{blue}{Welcome~To~OpenStudy}\]

OpenStudy (redheadangel):

the final value is normally rounded off so that there is no cents. Maybe it is $30,593, but all the answers shows some pennies. I am perplexed.......if you put the 5700 back on the answer36,293.09 which is the suggestion of ivettef365

OpenStudy (redheadangel):

5700 + 3650 =9350 + 537.56 = 9887.56 - [32862.44 ] you don't add the deductions

OpenStudy (redheadangel):

In questions like these all you have to do is take the gross income - IRA contributions - standard deduction -

OpenStudy (redheadangel):

5700 + 3650 + 537.56 + 918 + 1351.35 = 12,156.91. 42,750.00 - 12,156.91 = 30,593.09 ?<---taxable income

OpenStudy (anonymous):

Thank you :).